The length of the base of an isosceles triangle is 4 inches less than the length of one of the two equal sides of the triangles. If the perimeter is 32, find the three sides of the triangle.

Answer by checkley75(3666)

LET X=THE EQUAL SIDES.
THEN THE BASE=X-4.
THE PERIMETER IS THE SUM OF ALL THREE SIDES.
THUS WE HAVE THE FOLLOWING EQUATION FOR THE PERIMETER
X+X+(X-4)=32
3X-4=32
3X=32+4
3X=36
X=36/3
X=12 FOR THE EQUAL SIDES.
THUS THE BASE=12-4=8
PROOF
32=12+12+8
32=32